# Temperature-Noise Interplay in a Coupled Model of Opinion Dynamics

### v2

@@ Appendix A Limiting behavior $T\to 0$ for $q=2$
[…]
@@
[…]
@@ Appendix B Explicit formulas for $3\leq q\leq 6$ and the limiting behavior for $T\to\infty$ and $p=1$
The equation shows the critical value of noise for the base q-voter model. Requiring that $p_{3}^{*}(T_{c})=1$ leads to the following equation $01-2u^{3}-3u=$ 0, where $u=e^{-2/T_{c}}$ with only real solution $T=-2\ln^{-1}(r/2-1/r)$, where $r=[2(1+\sqrt{3})]^{1/3}$, i.e, $T_{c}\approx 1.7214$ – the critical temperature for monolayer q-Ising model. Thus, we see that the coupled model correctly gives limiting behavior of the base q-voter and q-Ising models.
@@
Here, similarly to $q=3$ case we can require $p_{3}^{*}(T_{c})=1$ which leads to $01-5u^{2}-12u=$ 0, where $u=e^{-4/T_{1}}$. The only real solution is $T_{1}=-4/\log\left[\frac{1}{5}\left(\sqrt{41}-6\right)\right]\approx 1.5886$, i.e., the value of the lower spinodal for the base q-Ising model.
@@
$m_{1,2}$ =±$=\pm\sqrt{\frac{B_{4}+\sqrt{B_{4}^{2}-4A_{4}C_{4}}}{2C_{4}}}$
$m_{3,4}$ =±$=\pm\sqrt{\frac{B_{4}-\sqrt{B_{4}^{2}-4A_{4}C_{4}}}{2C_{4}}}$
This allows us to calculate the value for the upper spinodal, making use of the fact that at this point $m_{1}$ and $m_{3}$ should be equal, which leads to $B_{4}^{2}-4A_{4}C_{4}=0$ and in turn to $05u^{4}-8u^{3}+10u^{2}-8u+1=$ 0 where $u=e^{-4/T_{2}}$. The only real, non-negative solution is then $T_{2}\approx 2.1092$.
@@
=$\underset{T\to\infty}{=}\frac{1}{5}$
As for $q=4$, the condition for the lower spinodal is $p_{5}^{*}(T_{1})=1$ and leads to $-3u^{5}-10u^{3}-10u+7=0$ where $u=e^{-2/T_{1}}$. The only real, non-negative solution is then $T_{1}\approx 3.1913$. To calculate the upper spinodal we once again use the fact that for higher terms ($D_{5}$, $E_{5}$, and $F_{5}$) disappear for $p=1$ and thus we can utlize Eq. (26), changing $A_{4}$, $B_{4}$ and $C_{4}$ coefficients into $A_{5}$, $B_{5}$ and $C_{5}$ which leads to
$04u^{10}-15u^{6}+20u^{5}-10u^{3}+1=$ 0 (30)
@@
=$\underset{T\to\infty}{=}\frac{5}{37}$
@@ Appendix C Landau approach for $q=5$
[…]
@@
Taking into account that $A_{5}^{L}$ has to be 0 at the critical and that $C^{L}_{5}(T,p_{5}^{*}(T))>0$ we conclude that the bistability region in the $q=5$ case disappears only when $T\to\infty$.

### v3

@@ Appendix A Limiting behavior $T\to 0$ for $q=2$
$\left\{.\begin{aligned} m_{0}&=0\\ m_{1,2}&=\pm\sqrt{\frac{5-7p}{1-p}}\\ m_{3,4}&=\pm 1\end{aligned}\right$. (20)
@@
where the stability of solutions has been resolved using the effective potential $V_{2}(m)=-\int F_{2}dm$ (see Fig. 8). The analysis proves that the solutions $m=\pm 1$ seen in Fig. 2b in the main text for $p>5/7$ are unstable. Additionally, we also see that the solution $m=\pm 1$ loses stability in favor of $m=\pm\sqrt{(5-7p)/(1-p)}$ once $p>2/3$.
@@ Appendix B Explicit formulas for $3\leq q\leq 6$ and the limiting behavior for $T\to\infty$ and $p=1$
The equation shows the critical value of noise for the base q-voter model. Requiring that $p_{3}^{*}(T_{c})=1$ leads to the following equation $1-2u^{3}-3u=0$, where $u=e^{-2/T_{c}}$ with only real solution $T=-2\ln^{-1}(r/2-1/r)$, where $r=[2(1+\sqrt{3})]^{1/3}$, i.e, $T_{c}\approx 1.7214$ – the critical temperature for monolayer q-Ising model. Thus, we see that the coupled model correctly gives limiting behavior of the base q-voter and q-Ising models.
@@
Here, similarly to $q=3$ case we can require $p_{3}^{*}(T_{c})=1$ which leads to $1-5u^{2}-12u=0$, where $u=e^{-4/T_{1}}$. The only real solution is $T_{1}=-4/\log\left[\frac{1}{5}\left(\sqrt{41}-6\right)\right]\approx 1.5886$, i.e., the value of the lower spinodal for the base q-Ising model.
@@
$m_{1,2}$ $=\pm\sqrt{\frac{B_{4}+\sqrt{B_{4}^{2}-4A_{4}C_{4}}}{2C_{4}}}$
$m_{3,4}$ $=\pm\sqrt{\frac{B_{4}-\sqrt{B_{4}^{2}-4A_{4}C_{4}}}{2C_{4}}}$
This allows us to calculate the value for the upper spinodal, making use of the fact that at this point $m_{1}$ and $m_{3}$ should be equal, which leads to $B_{4}^{2}-4A_{4}C_{4}=0$ and in turn to $5u^{4}-8u^{3}+10u^{2}-8u+1=0$ where $u=e^{-4/T_{2}}$. The only real, non-negative solution is then $T_{2}\approx 2.1092$.
@@
$\underset{T\to\infty}{=}\frac{1}{5}$
As for $q=4$, the condition for the lower spinodal is $p_{5}^{*}(T_{1})=1$ and leads to $3u^{5}-10u^{3}-10u+7=0$ where $u=e^{-2/T_{1}}$. The only real, non-negative solution is then $T_{1}\approx 3.1913$. To calculate the upper spinodal we once again use the fact that for higher terms ($D_{5}$, $E_{5}$, and $F_{5}$) disappear for $p=1$ and thus we can utlize Eq. (27), changing $A_{4}$, $B_{4}$ and $C_{4}$ coefficients into $A_{5}$, $B_{5}$ and $C_{5}$ which leads to
$4u^{10}-15u^{6}+20u^{5}-10u^{3}+1=0$ (31)
@@
$\underset{T\to\infty}{=}\frac{5}{37}$
@@ Appendix C Landau approach for $q=5$
The rate equation for $q=5$ and the relevant coefficients are given, respectively, by Eqs. 28 and 29. To follow the Landau approach, we calculate the effective potential $V_{5}=-\int F_{5}dm$ and expand it into a Taylor series around $m=0$, keeping the first three non-zero terms:
@@
Taking into account that $A_{5}^{L}$ has to be 0 at the critical line and that $C^{L}_{5}(T,p_{5}^{*}(T))>0$ we conclude that the bistability region in the $q=5$ case disappears only when $T\to\infty$.
Appendix D Analytical formula for the crtitical line $p^{*}_{q}(T)$
In this section, we derive a fully analytical formula for the critical line $p^{*}_{q}(T)$ given by Eq. (5) in terms of special functions (Gaussian hypergeometric functions), and then we prove that the limiting case $p^{*}_{q}(T\to\infty)$ coincides with the result obtained by Nyczka et al. for the critical value of noise in the monolayer q-voter model [9].
Let us start by restricting ourselves to only even values of $q$ – in this way, without losing the generality, we are able to facilitate further calculations. Using this assumption, we are able to notice that the form of $E_{qk}=\min\{1,\exp[2(q-2k)/T]\}$ implies immediately that $E_{ek}=1$ once $k\leq q/2$. This observation allows us to separate sums in the nominator and the denominator in Eq. (5) and rewrite it as
$p^{*}_{q}(T)=\frac{\sum\limits_{k=0}^{q/2}\tbinom{q}{k}\left(2q-2k-1\right)+\sum\limits_{k=q/2+1}^{q}\tbinom{q}{k}\left(2q-2k-1\right)E_{qk}}{\sum\limits_{k=0}^{q/2}\tbinom{q}{k}[(2^{q-1}-1)(1+2k-q)+q]+\sum\limits_{k=q/2+1}^{q}\tbinom{q}{k}[(2^{q-1}-1)(1+2k-q)+q]E_{qk}}.$ (39)
Let us now observe that the following summations hold true:
$\begin{split}\sum\limits_{k=0}^{q/2}\dbinom{q}{k}&=\frac{1}{2}\left[\binom{q}{\frac{q}{2}}+2^{q}\right]\\ \sum\limits_{k=0}^{q/2}\dbinom{q}{k}k&=2^{q-2}q\\ \sum\limits_{k=\frac{q}{2}+1}^{q}\dbinom{q}{k}E_{kq}&=e^{-4/T}\binom{q}{\frac{q}{2}+1}\,F1\\ \sum\limits_{k=\frac{q}{2}+1}^{q}\dbinom{q}{k}kE_{kq}&=\frac{1}{2}(q+2)e^{-4/T}\binom{q}{\frac{q}{2}+1}F1+\\ &+e^{-8/T}\binom{q}{\frac{q}{2}+2}\,F2\end{split}$ (40)
where, for brevity, we have used the following notation
$F1$ $={}_{2}F_{1}\left(1,1-\frac{q}{2};\frac{q}{2}+2;-e^{-4/T}\right)$ (41)
$F2$ $={}_{2}F_{1}\left(2,2-\frac{q}{2};\frac{q}{2}+3;-e^{-4/T}\right)$
to denote relevant Gaussian hypergeometric functions $_{2}F_{1}(a,b;c;z)$. The first two sums in Eq. (40) are straightforward and can be calculated by noticing that $\sum_{k=0}^{k=q}\binom{q}{k}=2^{q}$ and $\sum_{k=0}^{k=q}\binom{q}{k}k=2^{q-1}q$ and then making use of the fact that binomial coefficients are symmetrical.
The last two sums in Eq. (40) can be derived using the series expansion of the Gaussian hypergeometric function $_{2}F_{1}(a,b;c;z)=\sum_{k=0}^{\infty}\frac{z^{k}(a)_{k}(b)_{k}}{k!(c)_{k}}$ (see 15.1.1 in [43]), where $(x)_{k}$ is the rising factorial, i.e., $(x)_{k}=\prod_{k=1}^{n}(x-k+1)$. If $a$ or $b$ is a negative integer m-m, the infinite series reduces to a polynomial
$_{2}F_{1}(a,-m;c;z)=\sum_{k=0}^{m}\frac{z^{k}(-m)_{k}(b)_{k}}{k!(c)_{k}}$ (42)
(see 15.4.1 in [43] and note that $_{2}F_{1}(a,b;c;z)={}_{2}F_{1}(b,a;c;z)$). Taking into account relations Eq. (40) we are able to express $p^{*}_{q}(T)$ in a more concise form
$p^{*}_{q}(T)=\frac{4(q-3)e^{-4/T}\binom{q}{\frac{q}{2}+1}\,F1-8e^{-8/T}\binom{q}{\frac{q}{2}+2}\,F2+2\left((2q-1)\binom{q}{\frac{q}{2}}+2^{q}(q-1)\right)}{2\left(2q+3\ 2^{q}-6\right)e^{-4/T}\binom{q}{\frac{q}{2}+1}F1+4\left(2^{q}-2\right)e^{-8/T}\binom{q}{\frac{q}{2}+2}\,F2+\left(2^{q}-2-\left(2^{q}-4\right)q\right)\binom{q}{\frac{q}{2}}+2^{q}\left(2q+2^{q}-2\right)}$ (43)
Let us underline here that special functions used in this derivation are simply another way to express specific sums in Eq. (40). To obtain the limiting case $p^{*}_{q}(T\to\infty)$ let us note that we need to evaluate $\underset{T\to\infty}{F1}$ and $\underset{T\to\infty}{F2}$, i.e., $_{2}F_{1}\left(1,1-q/2;q/2+2;-1\right)$ and $_{2}F_{1}\left(1,2-q/2;q/2+3;-1\right)$. We shall use the following identities [43, 44]
$\,{}_{2}F_{1}(a,b;a-b+2;-1)$ $=\frac{\left(\sqrt{\pi}\Gamma(a-b+2)\right)\left(\frac{1}{\Gamma\left(\frac{a}{2}\right)\Gamma\left(\frac{a+3}{2}-b\right)}-\frac{1}{\Gamma\left(\frac{a+1}{2}\right)\Gamma\left(\frac{a}{2}-b+1\right)}\right)}{2^{a}(b-1)}$ (44)
$\,{}_{2}F_{1}(a,b;a-b+3;-1)$ $=\frac{\left(\sqrt{\pi}\Gamma(a-b+3)\right)\left(-\frac{a}{\Gamma\left(\frac{a}{2}+1\right)\Gamma\left(\frac{a+3}{2}-b\right)}+\frac{a(a+1)}{4\left(\Gamma\left(\frac{a+3}{2}\right)\Gamma\left(\frac{a}{2}-b+2\right)\right)}+\frac{1}{\Gamma\left(\frac{a+1}{2}\right)\Gamma\left(\frac{a}{2}-b+1\right)}\right)}{2^{a}((b-1)(b-2))}$
where $\Gamma(x)$ is the Euler gamma function, which allow us to write that
$\underset{T\to\infty}{F1}$ $=\left(\frac{1}{q}+\frac{1}{2}\right)\left(\frac{2^{q}}{\dbinom{q}{\frac{q}{2}}}-1\right)$ (45)
$\underset{T\to\infty}{F2}$ $=\frac{(q+2)(q+4)\left((q+2)\dbinom{q}{\frac{q}{2}}-2^{q+1}\right)}{4(q-2)q\dbinom{q}{\frac{q}{2}}}$
After substituting $F1$ and $F2$ in Eq. (43) with limiting values from Eq. (45) as well as taking the limit of $e^{-4/T}$ and $e^{-8/T}$ and some rudimentary algebra we finally arrive at
$p^{*}_{q}(T\to\infty)=\frac{2(q-1)}{2(q-1)+2^{q}}$ (46)
that matches the result obtained by Nyczka et al. [9].